# RecConv: Efficient Recursive Convolutions for Multi-Frequency Representations

### v1

@@ 4.2 Downstream Tasks
We evaluate RecNeXt’s transfer ablility on object detection and instance segmentation tasks. Following [34], we integrate RecNeXt into the Mask-RCNN framework [19] and conduct experiments on the MS-COCO 2017 dataset [35], presenting metrics of APbox and APmask. As shown in Table 4, RecNeXt demonstrates superior performance in terms of APbox and APmask while maintaining competitive latency. For instance, RecNeXt-M3 achieves a notable APbox of 41.7 and APmask of 38.6 while maintaining a latency of 5.2ms, offering an optimal trade-off between speed and precision. RecNeXt-M4 surpasses FastViT-SA24 by 1.5 APbox and 1.7 APmask with 4ms latency improvement, and increases the APbox and APmask by 0.6 compared to RepNext-M4 with only 0.5ms latency increase. However, RecNeXt-M5 encounters an performance saturation at 44.6 APbox and 40.6 APmask, albeit with a higher latency of 12.4ms compared to 9.3ms of RepNext-M5.
@@
Image Size Levels Kernel Size Receptive Field Params MACs Latency Throughput Top1 Upsample
(M) (G) (ms) (im/s) (%) Operation
$224\times 224$ $3\times 3$ [3, 3, 3, 3] 4.80 0.82 1.24 4479 74.64 Bilinear Interpolate
$7\times 7$ [7, 7, 7, 7] 4.96 0.87 1.27 3529 75.52
$11\times 1$ 11 [11, 11, 11, 11] 5.26 0.96 1.31 2591 75.41
@@
$5\times 5$ [80, 40, 20, 10] 5.16 0.87 1.36 384 76.03
$7\times 7$ [112, 56, 28, 14] 5.55 0.92 1.43 374 75.67
$384\times 384$ $3\times 3$ [3, 3, 3, 3] 4.80 2.40 2.07 1553 76.35
$7\times 7$ [7, 7, 7, 7] 4.96 2.55 2.20 1224 77.42
$11\times 1$ 11 [11, 11, 11, 11] 5.26 2.82 2.27 880 78.00
@@
$5\times 5$ [80, 40, 20, 10] 5.16 2.54 2.79 311 78.11
$7\times 7$ [112, 56, 28, 14] 5.55 2.69 2.84 295 78.03
$3\times 3$ [48, 24, 12, 6] 4.97 2.44 2.40 993 77.26 DWConv Transpose
$5\times 5$ [80, 40, 20, 10] 5.31 2.57 2.48 784 77.87
$7\times 7$ [112, 56, 28, 14] 5.81 2.75 2.63 576 78.24
@@
We perform semantic segmentation experiments on the ADE20K dataset [74], which consists of approximately 20K training images and 2K validation images across 150 categories. Intersection-over-Union (mIoU) is used as the evaluation metric. We followed the training protocol from previous works [32, 34] using the Semantic FPN framework [26]. As illustrated in Table 4, RecNeXt demonstrates favorable mIoU-latency trade-offs across various model sizes, with RecNeXt-M3 achieving the highest mIoU of 41.0% among other models of similar scale. RecNeXt-M4 and RecNeXt-M5 also achieve competitive mIoU and latency compared to the state-of-the-art models. These findings emphasize the critical role of efficient backbones with large receptive fields in advancing semantic segmentation, especially in computationally constrained environments.
@@ 4.3 Ablation Studies
Recursive decomposition. The integration of recursive decomposition expands the model’s receptive field (7→80) and improves multi-scale feature aggregation, leading to accuracy gains across resolutions. For $224\times 224$ images, decomposition with $5\times 5$ kernels achieves a top-1 accuracy of 76.03%, outperforming single-level configurations with larger kernel sizes of $7\times 7$ (75.52%) and $11\times 1$ 11 (75.41%). The benefits are even more pronounced for larger image resolutions ($384\times 384$), where the peak top-1 accuracy of 78.11% is achieved compared to 77.42%, without substantial changes in parameter count or FLOPs. However, the method increases latency (2.2ms $\rightarrow$ 2.8ms) and significantly reduces throughput (1200 img/s $\rightarrow$ 300 img/s). This partially due to the extensive employment of bilinear upsampling operation, which can be replaced by a transposed depthwise convolution for improved performance, albeit at the cost of increased computational complexity. In summary, the progressive decomposition allows the network to effectively aggregate multi-frequency representations while balancing computational complexity by distributing operations across different spatial scales.
@@
RecConv Variants. We present two RecConv variants (Figure 3 and 4) incorporating recursive decomposition along both spatial and channel dimensions. These variants improve computational efficiency but at the cost of accuracy. As summarized in Table 5, the variant in Figure 3 reduces parameter from 5.81M to 5.76M, FLOPs from 2.75G to 2.70G, and increases throughput from 576img/s to 636img/s, but results in a 0.33% drop in accuracy (from 78.24% to 77.91%). Meanwhile, the latency increases significantly due to the lack of adaptation of the transposed group convolution on the NPU. The variant in Figure 4 further reduces parameters by 0.4M and FLOPs by 0.1G, lowers latency by 0.3ms, and increases throughput by 200 img/s, but causes a more significant accuracy drop of 1.4%. This comparative analysis highlights the trade-off between upsampling strategies and decomposition methods in optimizing performance and computational efficiency. This study emphasizes design simplicity, using parameter-free bilinear upsampling operation and spatial-only decomposition to achieve a balanced solution.
Figure 3: RecConv Variant A. Recursive decomposition is implemented along both spatial and channel dimensions simultaneously. The shared downsample operation is replaced with group convolutions, where the number of groups matches the number of output channels. Bilinear upsample operations are substituted with transposed group convolutions, where the number of groups corresponds to the number of input channels.
Figure 4: RecConv Variant B. The recursive decomposition operation is jointly performed in both spatial and channel dimensions. The shared downsample operation is replaced with channel-wise split operations, followed by depthwise convolutions. Bilinear upsample operations are converted to transposed depthwise convolutions, and element-wise addition is substituted with channel-wise concatenation.
@@ 5 Conclusions
This paper introduces a straightforward and versatile recursive decomposition strategy that leverages small-kernel convolutions to construct multi-frequency representations, establishing a linear relationship between parameter growth and decomposition levels. This innovation guarantees that the computational complexity at each decomposition level follows a geometric progression, diminishing exponentially with increasing depth. This recursive design is compatible with various operations and can be extended to other modalities, though this study focuses on vision tasks using basic functions. Leveraging this approach, we construct RecConv as a plug-and-play module that can be seamlessly integrated into existing computer vision architectures. To the best of our knowledge, this is the first effective convolution kernel up to $80\times 80$ for resource-constrained vision tasks. Building on RecConv, we introduce RecNeXt, an efficient large-kernel vision backbone optimized towards resource-constrained scenarios. Extensive experiments across multiple vision benchmarks demonstrate RecNeXt’s superiority over current leading approaches without relying on structural reparameterization or neural architecture search.
@@
Limitations: While RecNeXt delivers competitive performance, its scalability lags behind the state-of-the-art models without knowledge distillation, where larger variants experience performance saturation and increased latency. Additionally, RecNeXt exhibits the lowest throughput among models of comparable parameter size due to the computational overhead of bilinear interpolation, which can be mitigated by employing transposed convolution. The recursive decomposition may introduce numerical instability during mixed precision training, which can be alleviated by using fixed-point or BFloat16 arithmetic. Compatibility issues with bilinear interpolation and transposed convolution on certain iOS versions may also result in performance degradation.

### v2

@@ 4.2 Downstream Tasks
We evaluate RecNeXt’s transfer ability on object detection and instance segmentation tasks. Following [38], we integrate RecNeXt into the Mask-RCNN framework [22] and conduct experiments on the MS-COCO 2017 dataset [39], presenting metrics of APbox and APmask. As shown in Table 4, RecNeXt demonstrates superior performance in terms of APbox and APmask while maintaining competitive latency. For instance, RecNeXt-M3 achieves a notable APbox of 41.7 and APmask of 38.6 while maintaining a latency of 5.2ms, offering an optimal trade-off between speed and precision. RecNeXt-M4 surpasses FastViT-SA24 by 1.5 APbox and 1.7 APmask with 4ms latency improvement, and increases the APbox and APmask by 0.6 compared to RepNext-M4 with only 0.5ms latency increase. However, RecNeXt-M5 encounters an performance saturation at 44.6 APbox and 40.6 APmask, albeit with a higher latency of 12.4ms compared to 9.3ms of RepNext-M5.
@@
Image Size Levels Kernel Size Receptive Field Params MACs Latency (ms) Throughput Top1 Upsample
(M) (G) NPU CPU (im/s) (%) Operation
$224\times 224$ $3\times 3$ [3, 3, 3, 3] 4.80 0.82 1.24 7.4 4479 74.64 -
$7\times 7$ [7, 7, 7, 7] 4.96 0.87 1.27 7.4 3529 75.52
$11\times 1$ 11 [11, 11, 11, 11] 5.26 0.96 1.31 7.5 2591 75.41
@@
$5\times 5$ [80, 40, 20, 10] 5.16 0.87 1.36 13.0 384 76.03
$7\times 7$ [112, 56, 28, 14] 5.55 0.92 1.43 13.0 374 75.67
@@
$7\times 7$ [112, 56, 28, 14] 5.35 0.89 - - 3221 75.48 Max Unpool
$384\times 384$ $3\times 3$ [3, 3, 3, 3] 4.80 2.40 2.07 13.2 1553 76.35 -
$7\times 7$ [7, 7, 7, 7] 4.96 2.55 2.20 13.2 1224 77.42
$11\times 1$ 11 [11, 11, 11, 11] 5.26 2.82 2.27 13.8 880 78.00
[…]
$5\times 5$ [80, 40, 20, 10] 5.16 2.54 2.79 20.6 311 78.11
$7\times 7$ [112, 56, 28, 14] 5.55 2.69 2.84 20.6 295 78.03
$3\times 3$ [48, 24, 12, 6] 4.97 2.44 2.40 41.5 993 77.26 DWConv Transpose
$5\times 5$ [80, 40, 20, 10] 5.31 2.57 2.48 63.5 784 77.87
$7\times 7$ [112, 56, 28, 14] 5.81 2.75 2.63 94.1 576 78.24
@@
We perform semantic segmentation experiments on the ADE20K dataset [80], which consists of approximately 20K training images and 2K validation images across 150 categories. Intersection-over-Union (mIoU) is used as the evaluation metric. We followed the training protocol from previous works [36, 38] using the Semantic FPN framework [30]. As illustrated in Table 4, RecNeXt demonstrates favorable mIoU-latency trade-offs across various model sizes, with RecNeXt-M3 achieving the highest mIoU of 41.0% among other models of similar scale. RecNeXt-M4 and RecNeXt-M5 also achieve competitive mIoU and latency compared to the state-of-the-art models. These findings emphasize the critical role of efficient backbones with large receptive fields in advancing semantic segmentation.
@@ 4.3 Ablation Studies
Recursive decomposition. The integration of recursive decomposition expands the model’s receptive field (7→80) and improves multi-scale feature aggregation, leading to accuracy gains across resolutions. For $224\times 224$ images, decomposition with $5\times 5$ kernels achieves a top-1 accuracy of 76.03%, outperforming single-level configurations with larger kernel sizes of $7\times 7$ (75.52%) and $11\times 1$ 11 (75.41%). The benefits are even more pronounced for larger image resolutions ($384\times 384$), where the peak top-1 accuracy of 78.11% is achieved compared to 77.42%, without substantial changes in parameter count or FLOPs. However, the method increases latency (2.2ms $\rightarrow$ 2.8ms) and significantly reduces throughput (1200 img/s $\rightarrow$ 300 img/s). This is partly due to the widespread employment of bilinear upsampling, which could be substituted with transposed depthwise convolution, nearest interpolation, or max unpooling, albeit at the expense of reduced accuracy. In summary, the progressive decomposition allows the network to effectively aggregate multi-frequency representations while balancing computational complexity by distributing operations across different spatial scales.
@@
RecConv Variants. We present RecConv variants (Figure 3) incorporating recursive decomposition along both spatial and channel dimensions. As summarized in Table 5, the variant with group convolutions reduces parameter from 5.81M to 5.76M, FLOPs from 2.75G to 2.70G, and increases throughput from 576 to 636, but results in a 0.33% drop in accuracy (from 78.24% to 77.91%). Meanwhile, the latency increases significantly due to the lack of adaptation of the transposed group convolution on the NPU. The variant using channel-wise concatenation further reduces parameters by 0.4M and FLOPs by 0.1G, lowers latency by 0.3ms, and increases throughput by 200, but causes a more significant accuracy drop of 1.4%.
Figure 3: RecConv Variant. Recursive decomposition is implemented along both spatial and channel dimensions. The shared downsampling layer is replaced by group convolutions, where the number of groups matches the number of output channels. Bilinear upsampling operations are substituted with transposed group convolutions, with groups corresponding to input channels.
Table 6: Ablation study over 300 epochs on ImageNet-1K at 256 image resolution with a batch size of 1024. Using two MLLA [18] variants as baselines (gray). Metrics include top-1 validation accuracy, NPU latency on an iPhone 13 running iOS 18 (mlmodel), CPU latency on an Intel(R) Xeon(R) Gold 6330 (onnx), throughput on a RTX4090 GPU, and peak mixed-precision training memory usage. $\dagger$” denotes the application of linear attention with RecConv. $\ast$” indicates partial CPU execution.
Upsample Params FLOPs Latency (ms) Throughput Memory Top1
[…]
- 3.6 0.9 - - 2601 48.2 76.3
Conv 4.1 0.9 34.7 26.9 2951 39.2 77.1
Interp 4.0 0.9 34.0 16.4 3196 38.7 77.1
[…]
- 13.9 3.1 - - 1126 97.3 82.3
Conv 14.5 2.9 72.6 53.8 1408 79.8 82.1
Interp 14.3 2.9 70.8 31.9 1545 78.6 81.9
[…]
RecConv Beyond. We apply RecConv to MLLA variants, replacing linear attention and downsampling layers. As shown in Table 6, modified models achieve comparable performance to baselines at similar FLOPs, with higher GPU throughput and reduced memory consumption. RecConv offers performance advantages in smaller models but underperformed the baseline as parameter count increases, while maintaining faster inference and lower memory usage. Models using nearest interpolation (Interp) for downsampling exhibit higher throughput and lower memory consumption compared to transposed depthwise convolution (Conv). Employing linear attention with RecConv also achieves comparable performance to the baseline model while reducing FLOPs, GPU latency, and training memory.
@@ 5 Conclusions
This paper introduces a simple yet flexible recursive decomposition strategy using small-kernel convolutions to construct multi-frequency representations, maintaining linear parameter growth with decomposition levels while ensuring computational complexity decreases exponentially (following geometric progression). Leveraging this framework, we construct RecConv as a plug-and-play module seamlessly integrable into existing vision architectures. To the best of our knowledge, this is the first attempt with effective receptive field up to $\mathbf{80\times 80}$ for resource-constrained vision tasks. Building on RecConv, we present RecNeXt, an efficient large-kernel vision backbone optimized towards resource-limited scenarios. We introduce two series of models: the “A” series uses linear attention and nearest interpolation, while the “M” series employs convolution and bilinear interpolation for simplicity and broader hardware compatibility (e.g., to address sub-optimal nearest interpolation support in some iOS versions). Extensive benchmarks demonstrate RecNeXt’s competitive performance over current leading approaches without requiring structural reparameterization or neural architecture search.
@@
Limitations: The “M” series exhibits the lowest throughput among models of comparable parameter size due to extensive use of bilinear interpolation (mitigable via nearest interpolation, max unpooling or transposed convolution). The recursive decomposition may introduce numerical instability during mixed precision training, which can be alleviated by using fixed-point or BFloat16 arithmetic.
Supplementary Material
6 RecConv with Linear Attention
Table 7: ImageNet-1K classification results. The “A” series uses linear attention and nearest interpolation with RecConv, while the “M” series employs convolution and bilinear interpolation for simplicity and broader hardware compatibility (e.g., to address suboptimal nearest interpolation support in some iOS versions). $\ast$” refers to hard knowledge distillation.
Params FLOPs Latency (ms) Throughput Top1
(M) (G) NPU CPU (im/s) (%)
M0 2.5 0.4 1.0 189 763 73.2 / 74.7$\ast$
A0 2.8 0.4 1.4 177 4902 73.6 / 75.0$\ast$
M1 5.2 0.9 1.4 361 384 78.0 / 79.2$\ast$
A1 5.9 0.9 1.9 334 2746 78.3 / 79.6$\ast$
M2 6.8 1.2 1.5 431 325 79.2 / 80.3$\ast$
A2 7.9 1.2 2.2 413 2327 79.6 / 80.8$\ast$
M3 8.2 1.4 1.6 482 314 79.6 / 80.9$\ast$
A3 9.0 1.4 2.4 447 2206 80.1 / 81.1$\ast$
M4 14.1 2.4 2.4 843 169 81.4 / 82.5$\ast$
A4 15.8 2.4 3.6 764 1265 81.6 / 82.5$\ast$
M5 22.9 4.7 3.4 1487 104 82.9 / 83.3$\ast$
A5 25.7 4.7 5.6 1376 721 83.1 / 83.5$\ast$
As detailed in Table 7, the RecNeXt models offer a compelling trade-off between accuracy, model size, and hardware-specific efficiency, achieving up to 83.1% (83.5% with knowledge distillation) Top-1 accuracy on ImageNet-1K with models from 2.5M to 25.7M parameters. The architecture provides two specialized paths: the M-series for accelerator-centric deployment with NPU latencies as low as 1.0ms, and the A-series for GPU-bound scenarios, where its linear attention and nearest interpolation design achieves higher throughput and better accuracy.
Table 8: Object detection and instance segmentation were evaluated using Mask R-CNN on MS-COCO 2017. Semantic segmentation results were obtained on ADE20K. Backbone FLOPs were measured with 512$\times$512 image resolution.
FLOPs (G) Object Detection Instance Segmentation Semantic
[…]
M3 7.3 41.7 63.4 45.4 38.6 60.5 41.4 41.0
A3 7.2 42.1 64.1 46.2 38.8 61.2 41.6 41.9
M4 12.4 43.5 64.9 47.7 39.7 62.1 42.4 43.6
A4 12.4 43.5 65.4 47.6 39.8 62.4 42.9 43.0
M5 24.6 44.6 66.3 49.0 40.6 63.5 43.5 46.0
A5 24.6 44.4 66.3 48.9 40.3 63.3 43.4 46.5
Table 8 presents a comparative analysis of our model families, evaluated across three distinct computational budgets on object detection, instance segmentation, and semantic segmentation tasks. At the most compact scale, the A3 model demonstrates superiority over M3 across all evaluated metrics while also being computationally efficient. As model complexity increases, a nuanced performance trade-off emerges. The M-series generally excels in object detection and instance segmentation. In contrast, the A-series consistently shows a competitive edge in semantic segmentation, with A3 achieving a notable +0.9 mIoU over M3 and A5 surpassing M5 by +0.5 mIoU.
7 Shape Bias Analysis
Figure 4: Shape bias comparison over 16 categories (modelvshuman). The vertical line is the average across categories.
We evaluate shape bias—the fraction of predictions based on shape rather than texture—using the modelvshuman benchmark [15] in Figure 4. A higher shape bias aligns with human perception and is thus more desirable. And we report quantitative results on the following datasets (Table 9): Cue-Conflict (CC), Sketch (SK), High-Pass (HP), Low-Pass (LP), Silhouette (SI), Power-Equalisation (PE) and EidolonI (EI). Our method outperforms other models across all 7 datasets while demonstrating superior shape bias.
Table 9: Model Performance on 7 OOD Datasets
Model FLOPs CC ($\uparrow$) SK ($\uparrow$) HP ($\uparrow$) LP ($\uparrow$) SI ($\uparrow$) PE ($\uparrow$) EI ($\uparrow$)
FastViT-T12 1.4 21.17 69.25 65.16 41.64 58.75 90.09 45.86
RepViT-M1.1 1.3 22.11 70.75 69.45 43.44 56.88 93.30 46.25
RepNeXt-M3 1.3 22.19 72.50 63.91 43.59 58.13 92.77 46.95
RecNeXt-M3 1.4 22.11 73.50 66.02 42.73 58.75 93.30 46.95
RecNeXt-A3 1.4 23.28 73.25 76.33 43.59 61.25 93.93 47.42
Figure 5: Grad-CAM on the MS-COCO validation dataset for RepViT-M1.1, FastViT-T12, RepNeXt-M3 and RecNeXt-M3.
8 More Comparisons
Table 10: ImageNet-1K classification performance. All models are trained for 300 epochs with a resolution of $224\times 224$. Where $\dagger$” denotes benchmark with Triton code. $\ast$” refers to hard knowledge distillation. NPU latency is measured on an iPhone 13 with models compiled by Core ML Tools. CPU latency is accessed on a Quad-core ARM Cortex-A57 processor.
Model Params FLOPs Latency (ms) Throughput Top1
(M) (G) NPU CPU (im/s) (%)
LSNet-T [71] 11.4 0.3 - 91 14708$\dagger$ 74.9 / 76.1$\ast$
RecNeXt-T 12.1 0.3 1.8 109 14181$\dagger$ / 13503 75.1 / 76.6$\ast$
LSNet-S [71] 16.1 0.5 - 139 9023$\dagger$ 77.8 / 79.0$\ast$
RecNeXt-S 15.8 0.7 2.0 188 8234$\dagger$ / 7902 78.3 / 79.6$\ast$
LSNet-B [71] 23.2 1.3 - 335 3996$\dagger$ 80.3 / 81.6$\ast$
RecNeXt-B 19.3 1.1 2.5 290 4385$\dagger$ / 4232 80.3 / 81.4$\ast$
We evaluate “T”, “S”, “B” variants on ImageNet-1K (Table 10), comparing them against the SOTA efficient architecture, LSNet [71]. The NPU latency is measured on an iPhone 13 with models compiled by Core ML Tools. The CPU latency is accessed on a Quad-core ARM Cortex-A57 processor in ONNX format. And the throughput is tested on a Nvidia RTX3090 with maximum power-of-two batch size that fits in memory. $\ast$” denotes hard knowledge distillation using the RegNetY-16GF [51] as the teacher model.
The smallest model, RecNeXt-T, achieves a Top-1 accuracy of 75.1%, which is a 0.2% improvement over LSNet-T, with a negligible increase in parameters and identical FLOPs. With knowledge distillation, RecNeXt-T reaches 76.6% accuracy, surpassing LSNet-T by 0.5%. The model demonstrates impressive hardware efficiency, with an NPU latency of 1.8ms, an ARM CPU latency of 109ms, and a GPU throughput of 14,181 images per second. RecNeXt-S achieves 78.3% Top-1 accuracy, outperforming LSNet-S by 0.5%, while having less parameters. The distilled version of RecNeXt-S widens this gap, reaching 79.6% accuracy. RecNeXt-B matches the 80.3% Top-1 accuracy of LSNet-B with fewer FLOPs and higher throughput. Across all model sizes, our architecture consistently delivers a competitive trade-off between accuracy, latency, and model size compared to LSNet. Table 11 presents the detailed architecture, where the “RecAttn Block” denotes partial channel operation [5, 76] of the RecConv block with linear attention.
Table 11: Architectural details of RecNeXt “T”, “S”, “B” variants.
Stage Resolution Type Config RecNeXt
T S B
stem $\frac{H}{2}\times\frac{W}{2}$ Convolution channels 16 32 32
$\frac{H}{4}\times\frac{W}{4}$ Convolution channels 32 64 64
$\frac{H}{8}\times\frac{W}{8}$ Convolution channels 64 128 128
1 $\frac{H}{8}\times\frac{W}{8}$ RecAttn Block channels 64 128 128
blocks 0 0 2
2 $\frac{H}{16}\times\frac{W}{16}$ RecAttn Block channels 128 256 256
blocks 2 2 8
3 $\frac{H}{32}\times\frac{W}{32}$ RecAttn Block channels 256 384 384
blocks 8 8 8
4 $\frac{H}{64}\times\frac{W}{64}$ RecAttn Block channels 512 512 512
blocks 10 10 12
9 Training Details
Table 12: Training recipe of “M” and “A” series and “T”, “S”, “B” variants of RecNeXt on ImageNet-1K. DropPath is set to $0.0$ for hard knowledge distillation.
Hyperparameters Config
optimizer AdamW
batch size $1024$
learning rate $32e-$ 3
LR schedule cosine
training epochs $300$
warmup epochs $5$
weight decay $0.025$
augmentation RandAug($9$, $0.5$)
color jitter $0.4$
gradient clip $0.02$
random erase $0.25$
label smooth $0.1$
mixup $0.8$
cutmix $1.0$
drop path M4/A4: $0.2$ M5/A5: $0.3$
T: $0.0$ S: $0.1$ B: $0.2$
For ImageNet-1K classification benchmark, we adopt the similar training recipe as RepViT [69] and LSNet [71]. All models are trained from scratch for 300 epochs on 224$\times$224 resolution. The initial learning rate is set to $32e-$ 3, and the total batch size is set to 1024. Table 12 provides the detailed training hyper-parameters.